For each line, use the given conditions to write an equation in the point-slope form and the slope-intercept form.
Slope = 4, passing through (1, 3)
Slope = 8, passing through (4, –1)
Slope = –5, passing through (–4, –2)
Slope = –2, passing through (0, –3)

Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

For each line, use the given conditions to write an equation in the point-slope form and the slope-intercept form.
Slope = 4, passing through (1,3)
Use
y-y1 = m*(x-x1) where m = slope and (x1,y1) is the given point.
y-3 = 4*(x-1)
y-3 = 4x-4
y = 4x-1
---------
Do the others like the one above.
--------------
Slope = 8, passing through (4,–1)
Slope = –5, passing through (–4,–2)
Slope = –2, passing through (0,–3)

Answer by solver91311(24713)



Use the point-slope form of the equation of a line:



where is the given slope, and is the given point.

For your first problem:



Then solve for to obtain the slope-intercept form. Do all four of them the same way.


John